1/7(p-3)=1/6(9p+5) Solve
Found 2 solutions by Alan3354, nerdybill:
Answer by Alan3354(69443)

1/7(p-3)=1/6(9p+5) Solve
-----------
Multiply by 42 (6*7)
6(p-3) = 7(9p+5)
6p-18 = 63p+35
57p = - 53
p = -53/57

Answer by nerdybill(7384)

1/7(p-3)=1/6(9p+5)
Multiply both sides by 42 (a common denominator) to get:
6(p-3)=7(9p+5)
6p-18=63p+35
-18=57p+35
-53 = 57p
-53/57 = p